OpenStudy (anonymous):

If the point (1, 5) lies on a circle whose center is at the origin, what is the equation of the circle?

OpenStudy (anonymous):

\[x^2+y^2 = r^2\]\[1^2+5^2=26\]\[x^2+y^2=26\]